OpenStudy (august899):

Find the arc length and area of a sector with radius r and central angle θ. r=5,θ=π/3.

OpenStudy (anonymous):

\[\frac{a}{r}=\theta\]solve for \(a\)

OpenStudy (anonymous):

Keep in mind that the definition of a radian is \[\frac{ arclength }{ radius }\]

OpenStudy (august899):

I thought I was supposed to do m/360*2π

OpenStudy (anonymous):

That's (arc lenght)